The heat generation rate in a plane wall of 0.24 m thickness is 0.4 MW/m^3. The wall is exposed on both sides to convection at 30°C.
a) Determine and compare the maximum temperatures for k=25, k=50, k=200 and k= 410 W/mK assuming h=250W/m^2 K.
b) Determine and compare the maximum temperatures for h=50, 250, 500 and 1000 W/m^2K with k =W/mK
Answer #1

t=0-24m Is • Igen agen = 0.4 W/m² Too = 30°c Too h As both sides conditions are same, then we can assume that both sides andDate: As we know maximum temperature should be surrounding because of Therefore, at X=0 or centre of wall same On both sides.As we can see from the calculatios that at constant convective heat transfer coeffiicient and increasing conductive heat transfer coefficient the temperature is increasing slowly while at constant conductive heat transfer coefficient and increasing convective heat transfer coefficient the temperature is decreasing at a very fast rate because of faster convection heat transfer.